Adults randomly selected for a poll were asked if they​ "favor or oppose using federal tax dollars to fund medical research using stem cells obtained from human​ embryos." Of the subjects​ surveyed, 308 were in​ favor, 294 were​ opposed, and 82 were unsure. A politician claims that people​ don't really understand the stem cell issue and their responses to such questions are random responses equivalent to a coin flip. Use a 0.01 significance level to test the claim that the proportion of subjects who respond in favor is equal to 0.5.

What is the p-value? (Round to three decimal places.)
Answer #1

H0: p = 0.5

Ha: p 0.5

Sample proportion = 308 / (308 + 294 + 82) = 308 / 684 = 0.4503

Test statistics

z = - p / sqrt( p (1 - p) / n)

= 0.4503 - 0.50 / sqrt( 0.50 * 0.50 / 684)  

= -2.60

This is test statistics value.

p-value = 2 * P( Z < z)

= 2 * P( Z < -2.60)

= 2 * 0.0047

= 0.009